OpenStudy (bbb911):

This is Geometry: To the nearest hundredth, find the surface area of a cone whose slant height is 9 centimeters and that has a base with radius 6 centimeters.

OpenStudy (anonymous):

169.65

OpenStudy (bbb911):

:D

OpenStudy (anonymous):

90*Pi for the whole cone including the base 90*Pi=282.74

OpenStudy (anonymous):

169.64 is without the base

OpenStudy (anonymous):

169.65*

OpenStudy (anonymous):

do you know the formula?

OpenStudy (bbb911):

not really

OpenStudy (anonymous):

ok I'll tell you: \[SA=\pi rl +\pi r^2\] r is the radius of the base and l is the slant length. The base portion of the formula is the Pi*r^2 which is just the area of a circle. The first area is the area of the cone part. To get this we cut the cone and unfold it so that it is a sector of a circle with radius l (slant length) and arc length 2*Pi*r. The portion of the whole circle is 2Pi*r/(2Pi*l) and the area of the whole circle would be Pi*l^2 so we multiply them together: \[\frac{2 \pi r}{2 \pi l}\pi l^2=\pi r l\] that's where the first part of the equation comes from

OpenStudy (bbb911):

Thank you soo much for that i will print it out :D

OpenStudy (anonymous):

you're welcome :)